Fertilizer with weed control costing $38 per 50- pound bag is to be mixed with a less expensive fertilizer costing $6.00 per 50- lb bag to make 16 bags of fertilizer that can be sold for $28.00 per bag. How many bags of cheaper fertilizer should be used.
Answer by stanbon(75887)

Fertilizer with weed control costing $38 per 50- pound bag is to be mixed with a less expensive fertilizer costing $6.00 per 50- lb bag to make 16 bags of fertilizer that can be sold for $28.00 per bag. How many bags of cheaper fertilizer should be used.
-------------------------------
Equation:
value + value = value
38x + 6(16-x) = 28*16
38x + 96 - 6x = 28*16
32x = 28*16 - 96
32x = 352
x = 11 (# of bags of $38 bags needed for the mix)
===========================
Cheers,
Stan H.